# Multi-Agent, Human-Agent and Beyond: A Survey on Cooperation in Social Dilemmas

### v1

@@ 1 Introduction
Social dilemmas (SDs, e.g., prisoner’s dilemma), spanning various domains including environmental pollution, public health crises, and resource management, present a fundamental conflict between personal interests and the common good (Nowak, 2006). While cooperation is beneficial for the collective, individuals are tempted to exploit or free-ride others’ efforts, potentially leading to a tragedy of the commons. Historically rooted in the study of biological altruism (Smith, 1982), the traditional research on cooperation in SDs has unveiled the pivotal roles of reciprocity and social preferences in fostering cooperative behaviors in human societies (Fehr et al., 2002; Rand and Nowak, 2013). Recently, propelled by advances in artificial intelligence (AI), this field has been undergoing a profound transformation—as AI agents now increasingly represent and engage with humans, our understanding of how cooperation emerges, evolves, and sustains in SDs is being significantly reshaped. This is particularly evident in two lines of research: multi-agent cooperation, where AI agents interact with each other in SDs, and human-agent cooperation, which examines the intricacies of human interactions with AI agents in SDs.
@@
A fundamental assumption of AI agents is their rationality. This raises a critical question: how can rational agents be steered towards effective cooperation, requiring them to overcome the lure of exploiting or free-riding others? This question has received much interest in recent studies of multi-agent cooperation, particularly in the context of sequential social dilemmas (SSDs) (Leibo et al., 2017). As opposed to normal-form SDs, SSDs are characterized by stochastic environments as well as larger state and policy spaces. It is shown that embedding human-like motives, such as fairness (Hughes et al., 2018) and social preferences (McKee et al., 2020), into agents’ rewards can promote cooperative behavior. In addition, mechanisms, like peer rewarding (Wang et al., 2018) and formal contracts (Christoffersen et al., 2023), have also shown potential in fostering cooperation among agents. On the other hand, it is shown that agents can safeguard themselves against other agents’ exploitation by pre-training a suite of policies and then choosing them adaptively in real-time (Lerer and Peysakhovich, 2017), or by influencing the future policies of their opponents (Foerster et al., 2018).
@@
In addition to multi-agent cooperation, human-agent cooperation is equally important, as interactions between humans and agents are becoming ubiquitous. This field is centered around two key questions. The first concerns the design of AI algorithms: how can agents cooperate with humans at a level comparable to human-human cooperation? Recent studies have shown that combining expert algorithms and reinforcement learning can lead to human-level cooperation in SDs involving both humans and agents (Crandall et al., 2018). Moreover, simpler strategies, such as the extortion and generous strategies, are also instrumental in enhancing human-agent cooperation (Hilbe et al., 2014). The second question delves into the socio-cognitive aspect: in what ways do human perceptions and reactions differ when dealing with AI agents compared to other humans? Interestingly, behavioral evidence has revealed that humans are prone to cooperate less and even exploit AI agents more when they are aware of the agents’ non-human nature (Ishowo-Oloko et al., 2019). To mitigate such human bias towards agents, one approach is to embed culturally relevant signals and emotionally expressive characteristics in AI agents (de Melo and Terada, 2019).
@@
Beyond multi-agent and human-agent cooperation, emerging research indicates that AI can enhance human-human cooperation in SDs, consequently, in turn, enriching the long-standing study of cooperative behaviors in human societies. Specifically, AI agents have been found to promote cooperation among humans by providing recommendations on partner selection (McKee et al., 2023). Moreover, analyses based on evolutionary game theory suggest that human-AI hybrid systems can surpass pure human populations in achieving higher cooperation levels (Guo et al., 2023).
@@
Each aforementioned field—multi-agent cooperation, human-agent cooperation, as well as the emergent field of leveraging AI agents to promote human-human cooperation—has attracted significant interest though, there lacks a comprehensive review that integrates insights from these fields. Existing reviews have touched on learning in SSDs, but those reviews primarily emphasize multi-agent reinforcement learning (MARL), with SSDs as specialized scenarios within mixed-motive MARL  (e.g., Zhang et al., 2021). On the other hand, numerous reviews have examined cooperative behaviors but predominantly within the context of human societies, and are restricted to normal-form SDs  (e.g., Rand and Nowak, 2013; Perc et al., 2017). In light of these observations, this survey aims to provide a comprehensive understanding that unifies these fields, and encapsulates the diverse facets of cooperation across these interconnected fields, as illustrated in Figure 1. This can not only reflect the current state of research at the intersection of AI and cooperation in SDs, but also illuminate potential paths for future investigations, particularly in the interplay of multi-agent, human-agent, and human-human cooperation.
@@ 2.1 Normal-Form Social Dilemmas
Social dilemmas (SDs) involve a conflict between immediate self-interest and longer-term collective interests (Van Lange et al., 2013). A normal-form SD is a general-sum normal-form game, where two players must choose one of two actions simultaneously: cooperation ($C$) or defection ($D$). Mutual cooperation yields the reward $R$ for both, mutual defection leads to the punishment payoff $P$ for both, and different choices give the cooperator the sucker’s payoff $S$ and the defector the temptation $T$. In an SD, these payoffs satisfy the following relationships: $R>P$ (mutual cooperation values higher than mutual defection), $R>S$ (mutual cooperation values higher than unilateral cooperation), $2R>T+S$ (social welfare of mutual cooperation values higher than that of unilateral cooperation), and either or both of the following two: $T>R$ (unilateral defection for greed values higher than mutual cooperation) or $P>S$ (unilateral defection for fear values higher than mutual cooperation). Table 1 shows three normal-form SDs.
Table 1: Payoff matrices of Prisoner’s Dilemma (PD), Stag Hunt (SH) and Hawk Dove (HD). The elements in the matrix are the payoffs for the row player, and the payoffs for the column player are at the transposed mirror position of the row player.
$\begin{array}[]{c|cc}\text{PD}&C&D\\ \hline\cr C&3&0\\ D&5&1\\ \end{array}$
$\begin{array}[]{c|cc}\text{SH}&C&D\\ \hline\cr C&5&0\\ D&3&1\\ \end{array}$
$\begin{array}[]{c|cc}\text{HD}&C&D\\ \hline\cr C&3&1\\ D&5&0\\ \end{array}$
@@ 2.2 Sequential Social Dilemmas
Sequential Social Dilemmas (SSDs) are social dilemmas happening over a sequence of time steps, defined on stochastic games that typically have larger state spaces and policy spaces than normal-form games and additionally consider the states of the decision-making environment. Formally, an $n$-agent SSD is a tuple $\langle\mathcal{M},\varPi=\varPi_{c}\bigsqcup\varPi_{d}\rangle$, where $\mathcal{M}$ is a stochastic game (also known as Markov game), $\varPi_{c}$ is a set of cooperative policies, and $\varPi_{d}$ is a set of defective policies. A stochastic game is a repeated game associated with probabilistic state transitions, where each agent chooses actions using a policy and all agents take actions simultaneously at the current environmental state; then, the environment will transfer to a new state, and each agent will receive a reward from the environment as a result of agents’ actions and the state transition. Consider a set $N_{c}$ of agents where each agent $j\in N_{c}$ adopts a policy $\pi^{j}\in\varPi_{c}$, and a set $N_{d}$ of agents where each agent $k\in N_{d}$ adopts a policy $\pi^{k}\in\varPi_{d}$, such that $N_{c}\cup N_{d}=N,N_{c}\cap N_{d}=\emptyset$. Denote the expected reward of any agent $i\in N$ by $R^{i}(l)$, with the number of agents adopting cooperative policies being $l=|N_{c}|$. For a SSD, the following conditions hold: $R^{j}(|N|)>R^{k}(0),\forall{j\in N_{c}}\ \forall{k\in N_{d}}$ (cooperators under full cooperation gain more than defectors under full defection); $R^{j}(|N|)>R^{j}(1),\forall{j\in N_{c}}$ (cooperators under full cooperation gain more than one cooperator under the others’ defection); and either or both of the following two: $R^{k}(l)>R^{j}(l),\forall{k\in N_{d}}\ \forall{j\in N_{c}}$ if $|N_{c}|>|N_{d}|$ (defectors gain more than cooperators when the number of cooperators is relatively large) or $R^{k}(l)>R^{j}(l),\forall{k\in N_{d}}\ \forall{j\in N_{c}}$ if $|N_{c}|<|N_{d}|$ (defectors gain more than cooperators when the number of cooperators is relatively small).
@@
Compared to the highly abstract normal-form SDs, SSDs have some characteristics that are more aligned with real-world scenarios, which are well-summarized by Leibo et al. (2017): First, real-world social dilemmas are temporally extended. Then, from an individual’s policy perspective, cooperation and defection are labels that apply to policies implementing individual strategic decisions, and one’s cooperativeness may be a graded quantity. Additionally, from the perspective of game participation, decisions to cooperate or defect occur only quasi-simultaneously, and sometimes decisions must be made when only having partial information about the state of the world and other players’ activities.
@@ 3 Multi-Agent Cooperation in SSDs
While multi-agent cooperation in sequential decision-making has been a long-standing topic in MARL that takes stochastic games as the standard model, this survey features a narrowed-down focus on multi-agent cooperation in SSDs. For multi-agent cooperation in stochastic games using MARL, we refer readers to (Busoniu et al., 2008) for a comprehensive survey.
Leibo et al. (2017) pioneered SSDs, defining the framework and proposing a modeling approach connecting two-agent SSDs and normal-form SDs through empirical game theoretic analysis (EGTA). Cooperative and defective policies are trained using independent deep reinforcement learning under various environmental settings. These policies are then evaluated to construct empirical game matrices. SSDs differ from normal-form SDs as agents in SSDs must learn both the policy and its effective execution, while normal-form SDs involve executing predefined actions without learning.
The challenges in solving SSDs include developing usable policies in stochastic games with large state and policy spaces, and addressing sub-optimal equilibria that independently maximizing the individual interest of each agent can lead to counterproductive outcomes in SSD. Fortunately, with the powerful generalization ability of deep neural networks, MARL offers a promising solution (Hernandez-Leal et al., 2019). This section reviews past works on solving SSDs with MARL on two lines as shown in Figure 1 (B). One line of research is concerned with how to promote multi-agent cooperation for higher social welfare by designing suitable reward functions through shaping intrinsic and external motivations. The other line focuses on finding good solutions to play against opponents with uncertain diverse policies, in order to exploit them or avoid being exploited by them.

### v2

@@ Abstract
keywords:
Social dilemma , Sequential social dilemma , Human-agent cooperation , Multi-agent reinforcement learning
journal: Neurocomputingequalequalfootnotetext: Equal contributions.
[…]
[inst1] organization=Northwestern Polytechnical University,city=Xi’an, postcode=710072, country=China
[…]
[inst2] organization=Tsinghua University,city=Beijing, postcode=100084, country=China
[…]
[inst3] organization=University of Auckland,city=Auckland, postcode=1001, country=New Zealand
[…]
[inst4] organization=Kyushu University,city=Fukuoka, postcode=816-8580, country=Japan
[…]
[inst5] organization=The Hong Kong Polytechnic University,city=Hong Kong, country=China
[…]
[inst6] organization=Shanghai Artificial Intelligence Laboratory,city=Shanghai, postcode=200030, country=China
@@ 1 Introduction
Social dilemmas (SDs, e.g., prisoner’s dilemma), spanning various domains including environmental pollution, public health crises, and resource management, present a fundamental conflict between personal interests and the common good [1]. While cooperation is beneficial for the collective, individuals are tempted to exploit or free-ride others’ efforts, potentially leading to a tragedy of the commons. Historically rooted in the study of biological altruism [2], the traditional research on cooperation in SDs has unveiled the pivotal roles of reciprocity, social preferences, and communications in fostering cooperative behaviors in human-human interactions [3, 4, 5, 6]. Recently, propelled by advances in artificial intelligence (AI), this field has been undergoing a profound transformation—as AI agents now increasingly represent and engage with humans, our understanding of how cooperation emerges, evolves, and sustains in SDs is being significantly reshaped. This is particularly evident in two lines of research: multi-agent cooperation, where AI agents interact with each other in SDs, and human-agent cooperation, which examines the intricacies of human interactions with AI agents in SDs.
@@
A fundamental assumption of AI agents is their rationality. This raises a critical question: how can rational agents be steered towards effective cooperation, requiring them to overcome the lure of exploiting or free-riding others? This question has received much interest in recent studies of multi-agent cooperation, particularly in the context of sequential social dilemmas (SSDs) [7]. As opposed to normal-form SDs, SSDs are characterized by stochastic environments as well as larger state and policy spaces. It is shown that embedding human-like motives, such as fairness [8] and social preferences [9], into agents’ rewards can promote cooperative behavior. In addition, mechanisms, like peer rewarding [10] and formal contracts [11], have also shown potential in fostering cooperation among agents. On the other hand, it is shown that agents can safeguard themselves against other agents’ exploitation by pre-training a suite of policies and then choosing them adaptively in real-time [12], or by influencing the future policies of their opponents [13].
@@
In addition to multi-agent cooperation, human-agent cooperation is equally important, as interactions between humans and agents are becoming ubiquitous. This field is centered around two key questions. The first concerns the design of AI algorithms: how can agents cooperate with humans at a level comparable to human-human cooperation? Recent studies have shown that combining expert algorithms and reinforcement learning can lead to human-level cooperation in SDs involving both humans and agents [14]. Moreover, simpler strategies, such as the extortion and generous strategies, are also instrumental in enhancing human-agent cooperation [15]. The second question delves into the socio-cognitive aspect: in what ways do human perceptions and reactions differ when dealing with AI agents compared to other humans? Interestingly, behavioral evidence has revealed that humans are prone to cooperate less and even exploit AI agents more when they are aware of the agents’ non-human nature [16]. To mitigate such human bias towards agents, one approach is to embed culturally relevant signals and emotionally expressive characteristics in AI agents [17].
@@
Beyond multi-agent and human-agent cooperation, emerging research indicates that AI can enhance human-human cooperation in SDs, consequently, in turn, enriching the long-standing study of cooperative behaviors in human societies. Specifically, AI agents have been found to promote cooperation among humans by providing recommendations on partner selection [18]. Moreover, analyses based on evolutionary game theory suggest that human-AI hybrid systems can surpass pure human populations in achieving higher cooperation levels [19].
@@
Each aforementioned field—multi-agent cooperation, human-agent cooperation, as well as the emergent field of leveraging AI agents to promote human-human cooperation—has attracted significant interest though, there lacks a comprehensive review that integrates insights from these fields. Existing reviews have touched on learning in SSDs, but those reviews primarily emphasize multi-agent reinforcement learning (MARL), with SSDs as specialized scenarios within mixed-motive MARL  [e.g., 20], without sufficiently addressing the close connection between SDs and human cooperation. On the other hand, numerous reviews have examined cooperative behaviors but predominantly within the context of human societies, and are restricted to normal-form SDs  [e.g., 5, 21], without fully acknowledging the increasingly significant role of AI in the study of SDs. In light of these observations, this survey aims to provide a comprehensive understanding that unifies these fields, and encapsulates the diverse facets of cooperation across these interconnected fields, as illustrated in Figure 1. This can not only reflect the current state of research at the intersection of AI and cooperation in SDs, but also illuminate potential paths for future investigations, particularly in the interplay of multi-agent, human-agent, and human-human cooperation. As AI becomes more integrated into human society, we call for increased research in the field of cooperative AI [22].
@@ 2.1 Normal-Form Social Dilemmas
Social dilemmas (SDs) involve a conflict between immediate self-interest and longer-term collective interests [23]. A normal-form SD is a general-sum normal-form game, where both players must choose one of two strategies simultaneously: cooperation ($C$) or defection ($D$). Mutual cooperation yields the reward $R$ for both, mutual defection leads to the punishment $P$ for both, and different choices give the cooperator the sucker’s payoff $S$ and the defector the temptation payoff $T$. In a SD, these payoffs satisfy the following relationships:
  • 1.
    $R>P$: mutual cooperation values higher than mutual defection and
  • 2.
    $R>S$: mutual cooperation values higher unilateral cooperation and
  • 3.
    $2R>T+S$: social welfare of mutual cooperation values higher than that of unilateral cooperation and
  • 4.
    $T>R$: unilateral defection for greed values higher than mutual cooperation or
  • 5.
    $P>S$: unilateral defection for fear values higher than mutual cooperation.
The utility functions of SD can be typically represented by the following payoff matrix $M$:
$\begin{array}[]{c|cc}&C&D\\ \hline\cr C&R&S\\ D&T&P\\ \end{array}.$
The elements in the matrix are the payoffs for the row player, and the payoffs for the column player are at the transposed mirror position of the row player. Next, we introduce three classic models of SDs: Prisoner’s Dilemma (PD), Stag Hunt (SH) and Hawk-Dove (HD).
Prisoner’s Dilemma
The PD is a classic social dilemma where cooperation results in the highest collective payoff, yet individuals acting in their own best interest tend to defect, leading to poorer outcomes for both involved players [24]. The payoff rankings for a PD specifically are $T>R>P>S$. An example of a payoff matrix for this scenario is presented below:
$\begin{array}[]{c|cc}\text{PD}&C&D\\ \hline\cr C&3&0\\ D&4&1\\ \end{array}.$
There is unique pure strategy Nash equilibrium  [^1] A Nash equilibrium in a game is a strategy profile from which no player can increase their payoff by unilaterally changing their strategy, provided the strategies of the other players remain unchanged [25]. $(D,D)$ in a PD.
Stag Hunt
In an SH, the highest collective payoff is achieved through mutual cooperation; however, players are tempted by lower-risk individual strategies due to fear of defection. The payoff structure adheres to the relationship $R>T\geq P>S$, indicating that defection may occur out of fear. Below is an illustrative payoff matrix for the SH:
$\begin{array}[]{c|cc}\text{SH}&C&D\\ \hline\cr C&5&0\\ D&2&1\\ \end{array}.$
There are two pure strategy Nash equilibrium $(C,C)$ and $(D,D)$ in an SH. Additionally, the SH also have a mixed strategy equilibrium, depending on the specific setup of the payoffs.
Hawk-Dove
The HD involves strategies of cooperation and defection in the context of resource competition among animals or individuals [25]. It reflects the trade-offs between risk and reward. The payoffs typically follow the order $T>R>S>P$, suggesting that players may choose aggression driven by greed. Below is an example of a payoff matrix for the HD:
$\begin{array}[]{c|cc}\text{HD}&C&D\\ \hline\cr C&3&1\\ D&4&0\\ \end{array}.$
There are two pure strategy Nash equilibrium $(C,D)$ and $(D,C)$ as well as a mixed strategy Nash equilibrium. Another two game models mathematically equivalent to the HD are Chicken game and Snowdrift game.
@@ 2.2 Sequential Social Dilemmas
Sequential Social Dilemmas (SSDs) are social dilemmas happening over a sequence of time steps and defined in stochastic games that consider the states of the decision-making environment. Formally, an $n$-agent SSD is a tuple $\langle\mathcal{M},\varPi=\varPi_{c}\bigsqcup\varPi_{d}\rangle$ [7], where $\mathcal{M}$ is a stochastic game (also known as Markov game), $\varPi_{c}$ is a set of cooperative policies, and $\varPi_{d}$ is a set of defective policies. A stochastic game is a repeated game associated with probabilistic state transitions, where each agent takes action based on its policy simultaneously at the current environmental state; then the environment will transform to a new state, and each agent will receive a reward from the environment as a result of agents’ actions and the state transition.
For an SSD, $\varPi_{c}$ and $\varPi_{d}$ induce the following manners. Consider a set $N_{c}$ of agents where each agent $j\in N_{c}$ adopts a policy $\pi^{j}\in\varPi_{c}$, and a set $N_{d}$ of agents where each agent $k\in N_{d}$ adopts a policy $\pi^{k}\in\varPi_{d}$, such that $N_{c}\cup N_{d}=N,N_{c}\cap N_{d}=\emptyset$. Denote the expected reward of the agent $i\in N$ by $R^{i}(l)$ in an episode of the stochastic game, with the number of agents adopting cooperative policies being $l=|N_{c}|$. Hughes et al. [8] propose conditions for an $n$-agent SSD to hold:
  • 1.
    $R^{j}(|N|)>R^{k}(0),\forall{j\in N_{c}}\ \forall{k\in N_{d}}$: cooperators under full cooperation gain more than defectors under full defection and
  • 2.
    $R^{j}(|N|)>R^{j}(1),\forall{j\in N_{c}}$: cooperators under full cooperation gain more than one cooperator under the others’ defection and
  • 3.
    $R^{k}(l)>R^{j}(l),\forall{k\in N_{d}}\ \forall{j\in N_{c}}$ if $|N_{c}|\geq|N_{d}|$: defectors gain more than cooperators when the number of cooperators is relatively large (for fear) or
  • 4.
    $R^{k}(l)>R^{j}(l),\forall{k\in N_{d}}\ \forall{j\in N_{c}}$ if $|N_{c}|\leq|N_{d}|$: defectors gain more than cooperators when the number of cooperators is relatively small (for greed).
@@
Compared to the highly abstract normal-form SDs, SSDs have some characteristics that are more aligned with real-world scenarios, which are well-summarized by Leibo et al. [7]: First, real-world social dilemmas are temporally extended. Then, from an individual’s policy perspective, cooperation and defection are labels that apply to policies implementing individual strategic decisions, and one’s cooperativeness may be a graded quantity. Additionally, from the perspective of game participation, decisions to cooperate or defect occur only quasi-simultaneously, and sometimes decisions must be made when only having partial information about the state of the world and other players’ activities.
Next we introduce several commonly used environments of SSDs.
Gathering [7]
Two agents aim to accumulate apples, represented by green pixels on the map. An agent earns a reward of 1 for each apple collected, after which the apple is temporarily removed and respawns after $N_{\text{apple}}$ frames. Agents are equipped with the capability to project a beam in alignment with their orientation. An agent that is struck by the beam twice is “tagged” and consequently removed from play for $N_{\text{tagged}}$ frames. Importantly, tagging does not confer rewards; its sole purpose may lie in reducing competition for apple resources. The illustration of Gathering is shown in Figure 2.
Figure 2: Illustration of Gathering [7].
Wolfpack [7]
Two agents (wolves) aim to chase a third agent (the prey). When either wolf touches the prey, all wolves within the capture radius receive a reward. The reward received by the capturing wolves is proportional to the number of wolves in the capture radius. The idea is that a lone wolf can capture the prey, but is at risk of losing the carcass to scavengers. However, when the two wolves capture the prey together, they can better protect the carcass from scavengers and hence receive a higher reward. A lone-wolf capture provides a reward of $r_{\text{lone}}$ and a capture involving both wolves is worth $r_{\text{team}}$. The illustration of Wolfpack is shown in Figure 3.
Figure 3: Illustration of Wolfpack [7].
Coins [12]
Two agents labeled in red and blue respectively, are tasked with picking up coins, also labeled red and blue respectively. If an agent picks up any coin by moving into the same position as the coin, it will earn a reward of 1. However, if the coin it pick up is in color of the other agent, the other agent receives a reward of -2. Thus, if both agents play greedily and pick up every coin, the expected reward for both agents is 0. The illustration of Coins is shown in Figure 4.
Figure 4: Illustration of Coins [12].
CleanUp [8]
The goal of all agents is to collect apples from a field. Each apple provides a reward of 1. The spawning of apples is controlled by a geographically separate aquifer that supplies water and nutrients. The river fills up with waste over time, lowering the respawn rate of apples linearly. For sufficiently high waste levels, no apples can spawn. To let apples to spawn, agents must clean waste in the river by firing clean beam. Furthermore, agents can also fire penalty beams, inflicting a reward of -50 reward on the hit agent at the cost of a -1 reward to themselves. The illustration of CleanUp is shown in Figure 5.
Figure 5: Illustration of CleanUp [8].
Harvest [8]
collecting apple provides a reward of 1. The apple regrowth rate varies across the map, dependent on the spatial configuration of uncollected apples: the more nearby apples, the higher the local regrowth rate. If all apples in a local area are harvested then none ever grow back until the game is reinitialized. Agents can fire penalty beams, inflicting a reward of -50 reward on the hit agent at the cost of a -1 reward to themselves. The illustration of Harvest is shown in Figure 6.
Figure 6: Illustration of Harvest [8].
@@ 3 Multi-Agent Cooperation in SSDs
While multi-agent cooperation in sequential decision-making has been a long-standing topic in MARL that takes stochastic games as the standard model, this survey features a narrowed-down focus on multi-agent cooperation in SSDs. For multi-agent cooperation in stochastic games using MARL, we refer readers to [26] for a comprehensive survey.
Leibo et al. [7] pioneered SSDs, defining the framework and proposing a modeling approach connecting two-agent SSDs and normal-form SDs through empirical game-theoretic analysis (EGTA). EGTA is a technique that employs agent-based simulation to generate data from game environments, thereby modeling game that are difficult to analyze [27]. In this work [7], cooperative and defective policies are trained using independent deep reinforcement learning under various environmental settings. These policies are then evaluated to construct empirical payoff matrices, which exhibit payoff relationships similar to that of normal-form SDs. However, from a learning perspective, SSDs differ from normal-form SDs as that agents in SSDs must learn effective execution of cooperation or defection rather than just learn to select between them.
The challenges in solving SSDs include developing usable policies in stochastic games with large state and policy spaces as mentioned above, as well as addressing sub-optimal equilibrium that independently maximizing the individual interest of each agent may lead to counterproductive outcomes in SSD. Fortunately, with the powerful generalization and representational ability of deep neural networks, MARL offers a promising solution to develop usable policies in high-dimensional stochastic games [28]. Therefore, we focus on addressing sub-optimal equilibrium in SSDs, which we call solve SSDs. This section reviews prevailing MARL-based works on solving SSDs on two lines as shown in Figure 1 (B). One line of research takes the planner’s view, which is concerned with how to promote multi-agent cooperation for higher social welfare by designing suitable reward functions through shaping intrinsic and external motivations.The other line adopts the participant’s view and tries to design agents to play against opponents with uncertain diverse policies, in order to exploit them or avoid being exploited by them. Their key difference lies in their perspectives on solving SSDs.